OpenStudy (anonymous):

Four polynomial functions are shown below: f(x) = x3 + 3x2 - 4x - 12 g(x) = x3 - 3x - 2 p(x) = x2 + 3x + 2 t(x) = x2 + x - 2 Which of the following functions has -1 and -2 as its zeros? f(x) g(x) p(x) t(x)

OpenStudy (solomonzelman):

factor each answer choice to find out.

OpenStudy (solomonzelman):

Don't know why you are posting a MATH question i n history, though :)

OpenStudy (solomonzelman):

\(\normalsize\color{black}{ f(x)=x^3 + 3x^2 - 4x - 12 }\) \(\normalsize\color{black}{ x^3 + 3x^2 - 4x - 12 =0}\) \(\normalsize\color{black}{ x^2(x+3) - 4(x+3) =0}\) \(\normalsize\color{black}{ (x^2-4)(x+3)=0}\) \(\normalsize\color{black}{ (x-2)(x+2)(x+3)=0}\) ZEROS, \(\normalsize\color{black}{ 2,~~-2,~~~~~3}\) .

OpenStudy (solomonzelman):

\(\normalsize\color{black}{ g(x)=x^3 - 3x - 2}\) will have non-integer zeros.

OpenStudy (solomonzelman):

\(\normalsize\color{black}{ p(x)=x^2 +3x+2}\) \(\normalsize\color{black}{ p(x)=(x+2)(x+1)}\) \(\normalsize\color{black}{ ~~-2,~~~~-1}\)